OpenStudy (anonymous):

Verify the equation below with each of the values listed for t to find a solution. 4 = 3t – 2

OpenStudy (muzzack):

well easy just isolate t then you've got 4+2=3t and then you have to divide both sides to take out the number multiplying so (3t)/3=(4+2)/3

OpenStudy (anonymous):

Im a bit tangled here im sorry. :(

OpenStudy (muzzack):

okay first isolate t so take everything around 3t

OpenStudy (muzzack):

to the other side

OpenStudy (muzzack):

so you have to take out -2 to the other side

OpenStudy (muzzack):

so you add both sides to cancel out the number on one side but the other you add 2

OpenStudy (muzzack):

try it

OpenStudy (anonymous):

thanks @Muzzack

OpenStudy (muzzack):

your welc :)